# A Step to Decouple Optimization in 3DGS

### v1

@@ C.3 Deformable Beta Splatting with AdamW-GS
Deformable Beta Splatting (DBS) [22] shares a similar pipeline design and same loss with 3DGSMCMC [10], as both are formulated under the Stochastic Gradient Langevin Dynamics framework. Unlike 3DGSMCMC, DBS replaces the Gaussian kernel with the Beta kernel and introduces the Spherical Beta function to better represent complex geometries and diverse appearance attributes. Consequently, the rendering process is reformulated as shown in Eq. 13.
Consistent with our experiments on 3DGSMCMC, we apply AdamW-GS to DBS, and the results are summarized in Table 9. Across eight of the nine scenes (excluding Treehill), we observe clear improvements in reconstruction quality when using AdamW-GS. For the Treehill scene, however, we find that DBS suffers from pronounced overfitting and triggers early stopping prematurely. Meanwhile, we observe that during the densification stage, has been early closed to zero, indicating that DAR cannot effectively exert its intended effect in this scenario. Our method may be less effective in such heavily overfitting scenario.
C.4 Vanilla 3DGS or 3DGSMCMC with AdamW-GS on OMMO Datasets
We provide results on the OMMO dataset [34], which contains large-scale outdoor scenes with long-range sequences. Our data processing follows the settings described in [10] and its associated Github repository 555https://github.com/ubc-vision/3dgs-mcmc. However, the image preprocessing failed for scene #10; thus, we report the average results for the left scenes. All results are summarized in Table 10. Our method demonstrates effectiveness on long sequence datasets as well.
@@ C.4 Vanilla 3DGS or 3DGSMCMC with AdamW-GS on OMMO Datasets
Appendix D Exploration Strategies
In Sec. 4.1, we observe that Sparse Adam exhibits less explosive characteristic. Part of this comes from the stability described in Observation 1, and part is attributable to the inherent bad transportation property of 3DGS [29], which we briefly discussed in Sec. 4.4. This phenomenon causes a larger portion of primitives to remain in a stable state during optimization with Sparse Adam—neither triggering further primitive generation nor enabling continued improvements in scene reconstruction. As a result, we obtain a primitive group with a smaller but inferior reconstruction quality. In Sec. 5 (Extra Exploration is Necessary), we examine the necessity of additional exploration through experiments that introduce noise regularization. Here, exploration is deliberately to encourage primitives to undergo larger movements, which is consistent with the notion of bad transportation described in [29]. We have given the discussion in Sec. 4.4
@@ Appendix D Exploration Strategies
In Sec. 5, we argue that the current AdamW-GS improves exploration. This enhancement arises from several factors: (1) RSR. We provide additional ablation studies showing that, when using only Sparse Adam together with RSR, training the vanilla 3DGS pipeline on Mip-NeRF 360 yields reconstruction quality comparable to vanilla 3DGS with Adam. The corresponding results are reported in Table 11 under “+ Only RSR (GS0)”. (2) The influence of DAR around saddle points. The analysis in [16] indicates that many primitives become trapped near saddle points, preventing further effective optimization. The presence of DAR encourages these primitives to continue participating in optimization rather than stagnating. (3) Improved gradient flow after removing redundant primitives. When a large number of redundant primitives are eliminated, the gradient flow within primitive groups becomes more coherent and effective. (4) Noise regularization for outdoor scenes. For outdoor datasets, we intentionally introduce noise regularization to further encourage extra exploration.
Existing work provides only limited discussion of exploration strategies for the 3DGS pipeline. Noise regularization, while helpful in certain cases, has inherent limitations and is not applicable to all datasets. Nevertheless, we argue that developing extra exploration strategies is a promising direction, as such mechanisms encourage primitives to explore a broader region of the space. Building on our current work, we introduce two additional exploration strategies: Adaptive Bound-Expanding Split [29] and Densification Extending. (For outdoor scenes, the following experiments still employ noise regularization; we further demonstrate that these strategies can be used jointly.)
@@
Densification Extending: We extend the original densification phase from 15,000 iterations to 25,000 iterations, leveraging the densification mechanism. Thanks to our the ability of our method to quickly penalize redundant primitives, this extension does not introduce the risk of memory explosion. The corresponding results are reported in Table 11 under “+ AdamW-GS (GS8) + Longer Densi”. Extending the densification phase further benefits indoor scenes and additionally reduces the number of active primitives.
Appendix E Failure Cases
@@
where is the depth of the primitive in the corresponding camera coordination.
Appendix F More about Related Work
@@
(1) Densification refinement: Several works optimize the densification rule [15, 7, 16]. SteepGS [16] mitigates redundancy through optimization-conditioned densification or alleviates sub-optimization by densification, but its reconstruction quality and pruning effectiveness remain inferior to learning-based methods. (2) Hand-crafted criterion: importance or redundancy criterion are designed to guide pruning after densification. LightGaussian [17] scores each Gaussian by opacity–transmittance–scale product times ray contributions. RadSplat [18] uses the maximum alpha–transmittance product across views. [11] propose the pipeline to prune primitives based on their degree of overlap. PUP-3DGS[19] proposes a principled sensitivity pruning score which is computed as a second-order approximation of the reconstruction error on the training views. (3) Learning-based pruning that leverages learning-based strategies to adaptively remove redundant primitives, often via mask. Compact3DGS [20] introduces Gaussian masks with L1 regularization on active mask count. MaskGaussion [12] models the existence of primitives, or measures the importance via learning. LP-3DGS[21] redesign the masking function to leverage the Gumbel-Sigmoid method. (4) Optimization-related operations: Optimization together with additional opacity L1 regularization has become a common technique for redundancy removal, even extending beyond Efficient 3DGS tasks [11, 20, 10, 22, 23].
Optimizing redundancy and training efficiency constitutes a subtopic of efficient 3DGS, corresponding to acceleration in inference and training (with partial overlap between the two). For inference, approaches include vector quantization [20], knowledge distillation [17], and entropy coding [35]. For training, efforts target backward acceleration and loss redesign [7]. These topics fall beyond the scope of this work: for fair comparison, we maintain the same loss function and refrain from using additional compression techniques.
@@
The development of adaptive first-order optimizers can be traced back to RProp [36]. AdaGrad [27] adapted the learning rate of features by estimated geometry and assign larger learning rate to infrequent features. RMSProp [37] further stabilized training by normalizing updates with an exponential moving average of squared gradients. Building on this, Adam [6] incorporated momentum into RMSProp through an exponential average of gradients, quickly becoming the default optimizer for modern DNNs [38]. Numerous refinements and extensions to Adam have since been proposed[39, 40]. Regularization has long been recognized as a key technique to improve model generalization and has also been employed as a tool to facilitate more effective optimization [25, 41, 42, 43, 44]. Importantly, [8] demonstrated that L2 regularization and weight decay are not equivalent, and further showed that weight decay provides a more appropriate formulation for Adam.
@@
Sparse Adam [7] introduces asynchronous updates that improve efficiency, yet the insufficient recognition of update-step coupling results in degraded performance. 3DGS-LM [45] proposes a tailored Levenberg-Marquardt optimizer for acceleration. However, it still relies on Adam for the initialization or densification. Although Adam has become a de facto indispensable optimizer for 3DGS, our understanding of its behavior in this context remains limited. A similar situation arises for regularization: despite the widespread use of various regularization techniques in 3DGS, there is currently little work examining the interplay between regularization and the optimizer, even though this has already been explored in the deep neural networks [8].
Appendix G Preparation for Decoupled Attribute Regularization
G.1 Hyperparameter Adjustment
@@
G.2 AdamW Style Decoupling
The AdamW-style decoupling implies an equal penalty applied to attributes, which can be formally expressed in Eq. 15 and Eq. 16. This formulation is equivalent to imposing a constant penalty on opacity [13]. Within the 3DGS-MCMC framework, we apply AdamW-style decoupling to opacity and evaluate it under three hyperparameter settings representing different strengths. To disentangle the effect of magnitude, we also experiment with constraining the magnitude of the AdamW-style decoupling, as given in Eq. 17. The results are summarized in Table 12. Consistent with the findings in Sec. G.1, excessively large hyperparameters hinder convergence, while overly small values render the regularization ineffective. A key limitation of AdamW-style decoupling is that it enforces uniform penalties across all primitives, despite their varying importance.
G.3 Additional discussion regarding the formulation in Eq.8
@@ G.2 AdamW Style Decoupling
G.4 Discussion Related to Implicit Update: Implicit Update Make Less Useful and Controllable
@@
Ineffective when stable: Implicit updates allow primitives in invisible viewpoints to be updated based on their current moment. Our reasoning, grounded in experimental results, is as follows. When primitives are in a stable state—more precisely, at saddle points, which are abundant in 3DGS when no external perturbation is applied [16]—implicit updates are largely ineffective while incurring additional computational cost. In vanilla 3DGS, we implemented AIU (see GS4 and GS3 in Sec. I), and observed negligible improvement.
@@
Less controllable when unstable: In contrast, when primitives are unstable, such as during the early training phase (e.g., densification, where newly generated primitives often have zero moment), or when the update step is relatively large, or when moment is rescaled, implicit updates can become problematic. We designed several experiments to highlight their uncontrollability. For instance, adopting AIU during densification in vanilla 3DGS with Sparse Adam can lead to more primitives but worse performance, as shown in GS6 in Sec. I, as even small update steps for invisible primitives introduce instability. Based on Obsidian 1, Sparse Adam tend to be stable, with which fewer primitives meet the gradient condition. AIU disrupts this stability: when primitives overfit certain views due to implicit updates, they induce larger gradients in other views, triggering unnecessary densification. The additional primitives do not necessarily improve rendering quality. In the Appendix Sec.I, we provide multiple AIU configurations: although in some cases the quality approaches that of vanilla 3DGS, the more common results are in worse quality, confirming the uncontrollable nature of implicit updates.
@@
Another related phenomenon is the dead primitives or more reallocated primitives (see Sec.5.1). The comparison between GS1 and GS3 in Table 1 (or can be found in Sec. I) shows that more dead primitives caused by Adam (0.232m vs. 0.048m). A natural conjecture is that when the update step, in a certain viewpoint, encourages primitives to become dead primitives, the implicit update accelerates this process, thereby producing more dead primitives. We provide some comparative experiments in Sec. I. (1) MC2 vs. MC3 in Sec. I, comparing the original 3DGSMCMC with Sparse Adam against the original 3DGSMCMC with Sparse Adam + AIU. (2) Similar comparisons can be found in GS3 and GS4, examining the effect of AIU on dead primitives when Sparse Adam is applied after densification ends. We observe that using AIU results in additional dead primitives. (3) Figure 3 likewise shows that adding AIU leads to more reallocated primitives during the densification stage of 3DGSMCMC. For certain scenes, such as Kitchen, the corresponding reconstruction quality decreases, further reflecting the increased number of dead primitives. This also indicates that the phenomenon is not part of a normal optimization process; in Sec. 5.1, we relate this behavior to over-effective regularization. (4) The comparison between MC6 and MC7 in Sec. I similarly shows that adding AIU after densification produces additional dead primitives as well; note that in our application RSR is only added during densification. (5) When training MaskGaussion with AdamW-GS, MaskGaussion will not suffer from the quality risk.
Appendix H Implementation Detail
H.1 AdamW-GS
@@
Since both opacity and scaling are strictly positive, the gradient of the L1 regularization reduces to a constant value of constant.
@@
H.2 Hyperparameter Selection
For vanilla 3DGS[1] and 3DGS-MCMC[10], all pipeline parameters follow the original settings, with noise parameters identical to those in 3DGS-MCMC. Operation for the opacity is similar as [13].
@@ H.2 Hyperparameter Selection
In general, the differences between these schedules arise from two main factors: (1) variations in the underlying pipelines (e.g., 3DGS vs. 3DGS-MCMC), which naturally require distinct schedules; and (2) variations across scenes, where different schedules are needed to obtain the best performance on indoor versus outdoor datasets. The discrepancies introduced by the pipelines themselves are unavoidable. Nevertheless, for each pipeline we design a conservative schedule—denoted StSSGS1 and StSSMC1—under which both 3DGS and 3DGS-MCMC outperform their original configurations. For outdoor scenes, which typically involve higher primitive generation or reallocation rates, we find that a more aggressive schedule further improves performance. All StSS schedules used in this paper are illustrated in Figure 9. Section I summarizes the results obtained by applying different StSS configurations across various pipelines and scenes.
@@
The three existing or designed operations provides guidance for selecting the parameters and in RSR. (1) The states of newly generated primitives in 3DGS, or reallocated primitives in 3DGS-MCMC, can be viewed as naturally constrained by an RSR with . (2) To properly control the magnitude of the update step , it is necessary to impose the relation . (3) A very small is consistent with the design philosophy of DAR, as we rely on such a small value to rescale the second moment and thereby activate regularization. Therefore, when and are chosen to be sufficiently small while satisfying , the configuration is safe. We conducted a search over different combinations of and . For larger configurations—for example, —RSR becomes less effective, and some combinations even lead to gradient explosion. For smaller values of and , the performance differences are minor, and we find that generally performs best. In Section I, we report results for both and , cited as MC19 and MC18. Hence, we adopt as default in all experiments.
@@
We use two DAR variants: opacity and scaling regularization. Regularization is constrained to the densification stage, as no clear benefit is observed in P-Op. This is consistent with the smaller max magnitude of adaptive gradients in P-Op (Figure 1). Opacity regularization starts after 3000 iteration, since a large percent of generated primitives initially have zero moment, motivating a delayed start.
For the two hyperparameters, we select values by considering activation derivatives, original settings, and Adam’s parameters. We set and, since scaling activations are two orders larger than opacity, . Comparative DAR experiments with and show that performs better (see experiments MC14 and MC17 in Sec. I). In our application, the only the most significant digit of is retained and then scaled down by one order of magnitude.
For opacity, is set to 10, guided by the max magnitude of adaptive gradients in Figure 1. Without , over-effective regularization occurs (see MC9 in Sec. I). Smaller values (e.g., 5 or 1) reduce reallocated primitives, leading to slight drops in PSNR and SSIM, though the MCMC framework still explores effectively. For scaling, we adopt by analogy with opacity, which yields good results; scaling is less sensitive to (e.g., shows little change in reconstruction quality; see MC8–MC17 in Sec. I).
Appendix I All Results

### v3

@@ E.3 Deformable Beta Splatting with AdamW-GS
Deformable Beta Splatting (DBS) (Liu et al., 2025a) shares a similar pipeline design and same loss with 3DGSMCMC (Kheradmand et al., 2024), as both are formulated under the Stochastic Gradient Langevin Dynamics framework. Unlike 3DGSMCMC, DBS replaces the Gaussian kernel with the Beta kernel and introduces the Spherical Beta function to better represent complex geometries and diverse appearance attributes. Consequently, the rendering process is reformulated as shown in Eq. 13.
Consistent with our experiments on 3DGSMCMC, we apply AdamW-GS to DBS, and the results are summarized in Table 9. Across eight of the nine scenes (excluding Treehill), we observe clear improvements in reconstruction quality when using AdamW-GS. For the Treehill scene, however, we find that DBS suffers from pronounced overfitting and triggers early stopping prematurely. Meanwhile, we observe that during the densification stage, has been closed early to zero, indicating that DAR cannot effectively exert its intended effect in this scenario. Our method may be less effective in such heavily overfitting scenario.
E.4 Vanilla 3DGS or 3DGSMCMC with AdamW-GS on OMMO Datasets
We provide results on the OMMO dataset (Lu et al., 2023), which contains large-scale outdoor scenes with long-range sequences. Our data processing follows the settings described in (Kheradmand et al., 2024) and its associated Github repository 555https://github.com/ubc-vision/3dgs-mcmc. However, the image preprocessing failed for scene #10; thus, we report the average results for the left scenes. All results are summarized in Table 10. Our method demonstrates effectiveness on long sequence datasets as well.
@@ E.4 Vanilla 3DGS or 3DGSMCMC with AdamW-GS on OMMO Datasets
Appendix F Exploration Strategies
In Sec. 4.1, we observe that Sparse Adam exhibits less explosive characteristic. Part of this comes from the stability described in Observation 1, and part is attributable to the inherent bad transportation property of 3DGS (Jung et al., 2024), which we briefly discussed in Sec. 4.4. This phenomenon causes a larger portion of primitives to remain in a stable state during optimization with Sparse Adam—neither triggering further primitive generation nor enabling continued improvements in scene reconstruction. As a result, we obtain a primitive group with a smaller but inferior reconstruction quality. In Sec. 5 (Extra Exploration is Necessary), we examine the necessity of additional exploration through experiments that introduce noise regularization. Here, exploration is deliberately to encourage primitives to undergo larger movements, which is consistent with the notion of bad transportation described in (Jung et al., 2024). We have given the discussion in Sec. 4.4.
@@ Appendix F Exploration Strategies
In Sec. 5, we argue that the current AdamW-GS improves exploration. This enhancement arises from several factors: (1) RSR. We provide additional ablation studies showing that, when using only Sparse Adam together with RSR, training the vanilla 3DGS pipeline on Mip-NeRF 360 yields reconstruction quality comparable to vanilla 3DGS with Adam. The corresponding results are reported in Table 11 under “+ Only RSR (GS0)”. (2) The influence of DAR around saddle points. The analysis in (Wang et al., 2025) indicates that many primitives become trapped near saddle points, preventing further effective optimization. The presence of DAR encourages these primitives to continue participating in optimization rather than stagnating. (3) Improved gradient flow after removing redundant primitives. When a large number of redundant primitives are eliminated, the gradient flow within primitive groups becomes more coherent and effective. (4) Noise regularization for outdoor scenes. For outdoor datasets, we intentionally introduce noise regularization to further encourage extra exploration.
Existing work provides only limited discussion of exploration strategies for the 3DGS pipeline. Noise regularization, while helpful in certain cases, has inherent limitations and is not applicable to all datasets. Nevertheless, we argue that developing extra exploration strategies is a promising direction, as such mechanisms encourage primitives to explore a broader region of the space. Building on our current work, we introduce two additional exploration strategies: Adaptive Bound-Expanding Split (Jung et al., 2024) and Densification Extending. (For outdoor scenes, the following experiments still employ noise regularization; we further demonstrate that these strategies can be used jointly.)
@@
Densification Extending: We extend the original densification phase from 15,000 iterations to 25,000 iterations, leveraging the densification mechanism. Thanks to the ability of our method to quickly penalize redundant primitives, this extension does not introduce the risk of memory explosion. The corresponding results are reported in Table 11 under “+ AdamW-GS (GS8) + Longer Densi”. Extending the densification phase further benefits indoor scenes and additionally reduces the number of active primitives.
Appendix G Failure Cases
@@
where is the depth of the primitive in the corresponding camera coordinations.
Appendix H More about Related Work
@@
(1) Densification refinement: Several works optimize the densification rule (Fang & Wang, 2024; Mallick et al., 2024; Wang et al., 2025). SteepGS (Wang et al., 2025) mitigates redundancy through optimization-conditioned densification or alleviates sub-optimization by densification, but its reconstruction quality and pruning effectiveness remain inferior to learning-based methods. (2) Hand-crafted criterion: importance or redundancy criteria are designed to guide pruning after densification. LightGaussian (Fan et al., 2024) scores each Gaussian by opacity–transmittance–scale product times ray contributions. RadSplat (Niemeyer et al., 2025) uses the maximum alpha–transmittance product across views. (Papantonakis et al., 2024) propose the pipeline to prune primitives based on their degree of overlap. PUP-3DGS(Hanson et al., 2025) proposes a principled sensitivity pruning score which is computed as a second-order approximation of the reconstruction error on the training views. (3) Learning-based pruning that leverages learning-based strategies to adaptively remove redundant primitives, often via mask. Compact3DGS (Lee et al., 2024) introduces Gaussian masks with L1 regularization on active mask count. MaskGaussian (Liu et al., 2025b) models the existence of primitives, or measures the importance via learning. LP-3DGS(Zhang et al., 2024a) redesigns the masking function to leverage the Gumbel-Sigmoid method. (4) Optimization-related operations: Optimization together with additional opacity L1 regularization has become a common technique for redundancy removal, even extending beyond Efficient 3DGS tasks (Papantonakis et al., 2024; Lee et al., 2024; Kheradmand et al., 2024; Liu et al., 2025a; Svitov et al., 2024).
Optimizing redundancy and training efficiency constitute a subtopic of efficient 3DGS, corresponding to acceleration in inference and training (with partial overlap between the two). For inference, approaches include vector quantization (Lee et al., 2024), knowledge distillation (Fan et al., 2024), and entropy coding (Huang et al., 2025). For training, efforts target backward acceleration and loss redesign (Mallick et al., 2024). These topics fall beyond the scope of this work: for fair comparison, we maintain the same loss function and refrain from using additional compression techniques.
@@
The development of adaptive first-order optimizers can be traced back to RProp (Riedmiller & Braun, 1993). AdaGrad (Duchi et al., 2011) adapted the learning rate of features by estimated geometry and assigns larger learning rate to infrequent features. RMSProp (Hinton et al., 2012) further stabilized training by normalizing updates with an exponential moving average of squared gradients. Building on this, Adam (Kinga et al., 2015) incorporated momentum into RMSProp through an exponential average of gradients, quickly becoming the default optimizer for modern DNNs (Vaswani et al., 2017). Numerous refinements and extensions to Adam have since been proposed(You et al., 2019; Chen et al., 2023). Regularization has long been recognized as a key technique to improve model generalization and has also been employed as a tool to facilitate more effective optimization (Srivastava et al., 2014; Shalev-Shwartz & Ben-David, 2014; Andriushchenko et al., 2023; Brown et al., 2020; Radford et al., 2021). Importantly, Loshchilov & Hutter (2017) demonstrated that L2 regularization and weight decay are not equivalent, and further showed that weight decay provides a more appropriate formulation for Adam.
@@
Sparse Adam (Mallick et al., 2024) introduces asynchronous updates that improve efficiency, yet the insufficient recognition of update-step coupling results in degraded performance. 3DGS-LM (Höllein et al., 2025) proposes a tailored Levenberg-Marquardt optimizer for acceleration. However, it still relies on Adam for the initialization or densification. Although Adam has become a de facto indispensable optimizer for 3DGS, our understanding of its behavior in this context remains limited. A similar situation arises for regularization: despite the widespread use of various regularization techniques in 3DGS, there is currently little work examining the interplay between regularization and the optimizer, even though this has already been explored in the deep neural networks (Loshchilov & Hutter, 2017).
Appendix I Preparation for Decoupled Attribute Regularization
I.1 Hyperparameter Adjustment
@@
I.2 AdamW Style Decoupling
The AdamW-style decoupling implies an equal penalty applied to attributes, which can be formally expressed in Eq. 15 and Eq. 16. This formulation is equivalent to imposing a constant penalty on opacity (Rota Bulò et al., 2025). Within the 3DGS-MCMC framework, we apply AdamW-style decoupling to opacity and evaluate it under three hyperparameter settings representing different strengths. To disentangle the effect of magnitude, we also experiment with constraining the magnitude of the AdamW-style decoupling, as given in Eq. 17. The results are summarized in Table 12. Consistent with the findings in Appendix Sec. I.1, excessively large hyperparameters hinder convergence, while overly small values render the regularization ineffective. A key limitation of AdamW-style decoupling is that it enforces uniform penalties across all primitives, despite their varying importance.
I.3 Additional discussion regarding the formulation in Eq.8
@@ I.2 AdamW Style Decoupling
I.4 Discussion Related to Implicit Update
@@
Ineffective when stable: Implicit updates allow primitives in invisible viewpoints to be updated based on their current moment. Our reasoning, grounded in experimental results, is as follows. When primitives are in a stable state—more precisely, at saddle points, which are abundant in 3DGS when no external perturbation is applied (Wang et al., 2025)—implicit updates are largely ineffective while incurring additional computational cost. In vanilla 3DGS, we implemented AIU (see GS4 and GS3 in Sec. K), and observed negligible improvement.
@@
Less controllable when unstable: In contrast, when primitives are unstable, such as during the early training phase (e.g., densification, where newly generated primitives often have zero moment), or when the update step is relatively large, or when moment is rescaled, implicit updates can become problematic. We designed several experiments to highlight their uncontrollability. For instance, adopting AIU during densification in vanilla 3DGS with Sparse Adam can lead to more primitives but worse performance, as shown in GS6 in Appendix Sec. K, as even small update steps for invisible primitives introduce instability. Based on Obsidian 1, Sparse Adam tends to be stable, with which fewer primitives meet the gradient condition. AIU disrupts this stability: when primitives overfit certain views due to implicit updates, they induce larger gradients in other views, triggering unnecessary densification. The additional primitives do not necessarily improve rendering quality. In the Appendix Sec.K, we provide multiple AIU configurations: although in some cases the quality approaches that of vanilla 3DGS, the more common results are in worse quality, confirming the uncontrollable nature of implicit updates.
@@
Another related phenomenon is the dead primitives or more reallocated primitives (see Sec.5.1). The comparison between GS1 and GS3 in Table 1 (or can be found in Appendix Sec. K) shows that more dead primitives are caused by Adam (0.232m vs. 0.048m). A natural conjecture is that when the update step, in a certain viewpoint, encourages primitives to become dead primitives, the implicit update accelerates this process, thereby producing more dead primitives. We provide some comparative experiments in Appendix Sec. K. (1) MC2 vs. MC3 in Appendix Sec. K, comparing the original 3DGSMCMC with Sparse Adam against the original 3DGSMCMC with Sparse Adam + AIU. (2) Similar comparisons can be found in GS3 and GS4, examining the effect of AIU on dead primitives when Sparse Adam is applied after densification ends. We observe that using AIU results in additional dead primitives. (3) Figure 3 likewise shows that adding AIU leads to more reallocated primitives during the densification stage of 3DGSMCMC. For certain scenes, such as Kitchen, the corresponding reconstruction quality decreases, further reflecting the increased number of dead primitives. This also indicates that the phenomenon is not part of a normal optimization process; in Sec. 5.1, we relate this behavior to over-effective regularization. (4) The comparison between MC6 and MC7 in Appendix Sec. K similarly shows that adding AIU after densification produces additional dead primitives as well; note that in our application RSR is only added during densification. (5) When training MaskGaussian with AdamW-GS, MaskGaussian will not suffer from the quality risk.
Appendix J Implementation Details
J.1 AdamW-GS
@@
Since both opacity and scaling are strictly positive, the gradient of the L1 regularization reduces to a constant value of .
@@
J.2 Hyperparameter Selection
For vanilla 3DGS(Kerbl et al., 2023) and 3DGS-MCMC(Kheradmand et al., 2024), all pipeline parameters follow the original settings, with noise parameters identical to those in 3DGS-MCMC. Operation for the opacity is similar as (Rota Bulò et al., 2025).
@@
In general, the differences between these schedules arise from two main factors: (1) variations in the underlying pipelines (e.g., 3DGS vs. 3DGS-MCMC), which naturally require distinct schedules; and (2) variations across scenes, where different schedules are needed to obtain the best performance on indoor versus outdoor datasets. The discrepancies introduced by the pipelines themselves are unavoidable. Nevertheless, for each pipeline we design a conservative schedule—denoted StSSGS1 and StSSMC1—under which both 3DGS and 3DGS-MCMC outperform their original configurations. For outdoor scenes, which typically involve higher primitive generation or reallocation rates, we find that a more aggressive schedule further improves performance. All StSS schedules used in this paper are illustrated in Figure 9. Appendix Sec.K summarizes the results obtained by applying different StSS configurations across various pipelines and scenes.
@@
The three existing or designed operations provide guidance for selecting the parameters and in RSR. (1) The states of newly generated primitives in 3DGS, or reallocated primitives in 3DGS-MCMC, can be viewed as naturally constrained by an RSR with . (2) To properly control the magnitude of the update step , it is necessary to impose the relation . (3) A very small is consistent with the design philosophy of DAR, as we rely on such a small value to rescale the second moment and thereby activate regularization. Therefore, when and are chosen to be sufficiently small while satisfying , the configuration is safe. We conducted a search over different combinations of and . For larger configurations—for example, —RSR becomes less effective, and some combinations even lead to gradient explosion. For smaller values of and , the performance differences are minor, and we find that generally performs best. In Appendix Sec.K, we report results for both and , cited as MC19 and MC18. Hence, we adopt as default in all experiments.
@@
We use two DAR variants: opacity and scaling regularization. Regularization is constrained to the densification stage, as no clear benefit is observed in P-Op. This is consistent with the smaller max magnitude of adaptive gradients in P-Op (Figure 1). Opacity regularization starts after 3000 iterations, since a large percent of generated primitives initially have zero moment, motivating a delayed start.
For the two hyperparameters, we select values by considering activation derivatives, original settings, and Adam’s parameters. We set and, since scaling activations are two orders larger than opacity, . Comparative DAR experiments with and show that performs better (see experiments MC14 and MC17 in Appendix Sec. K). In our application, only the most significant digit of is retained and then scaled down by one order of magnitude.
For opacity, is set to 10, guided by the max magnitude of adaptive gradients in Figure 1. Without , over-effective regularization occurs (see MC9 in Appendix Sec. K). Smaller values (e.g., 5 or 1) reduce reallocated primitives, leading to slight drops in PSNR and SSIM, though the MCMC framework still explores effectively. For scaling, we adopt by analogy with opacity, which yields good results; scaling is less sensitive to (e.g., shows little change in reconstruction quality; see MC8–MC17 in Appendix Sec. K).
Appendix K All Results